OpenStudy (anonymous):

Find the slope of the line tangent to p(x)=(2^x)/(x^2+1) at x=3

OpenStudy (anonymous):

\[p(x)=\frac{ 2^x }{ x^2+1 }\] is the equatio from teh problem

OpenStudy (campbell_st):

find the 1st derivative and substitute x = 3 to find the slope.

OpenStudy (anonymous):

what is the 1st derivative I can't figure it out

OpenStudy (anonymous):

Quotient rule with a bit of a trick to it. The derivative of any constant c^x is c^x * ln(c). Something nice to remember and we will have to use it here. So let's review the quotient rule: \[\frac{ g'(x)h(x) - g(x)h'(x)}{ h(x)^2}\] Where g(x) is the function in the numerator, h(x) the function in the denominator. So apply this to our function: The derivative of the top: \[\frac{ d }{ dx } (2^x) = 2^x * \ln(2)\] Derivative of the bottom: \[\frac{ d }{ dx } (x^2 + 1) = 2x\] (Constants equal zero) Plug it all in, square the denominator: \[\frac{ (2^x \ln(x))(x^2 + 1) - 2^x * 2x }{ (x^2 + 1)^2 }\] Simplify if you please. If I can also get someone to check my work that would be wonderful.

OpenStudy (campbell_st):

oops you final should read (\[\frac{dy}{dx} = \frac{2^xlog(2)\times(x^2 +1) - 2^x \times 2x}{(x^2 + 1)^2}\]

OpenStudy (anonymous):

Where was my mistake?

OpenStudy (anonymous):

the final answer is .7044 correct?

OpenStudy (campbell_st):

you have \[2^xlog(x).... \] and shouldn't it be \[2^x \log(2)\]

OpenStudy (anonymous):

is my answer correct?

OpenStudy (anonymous):

Oh yes, didn't catch that. Thanks for the correction. And hold on farmergirl I'm checking it

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

here is the wolfram alpha version... http://www.wolframalpha.com/input/?i=differentiate+y+%3D+%282%5Ex%29%2F%28x%5E2+%2B1%29

OpenStudy (anonymous):

This derivative is correct according to Wolfram Alpha: \[ \frac{(2^x*\ln(2))(x^2+1) − 2^x*2x }{ (x^2+1)^2 }\] And no farmergirl, your answer is incorrect. This is a bit of a tedious derivative to computate so if you give up, where's the link to the answer : https://www.wolframalpha.com/input/?i=%28%282^x*ln%282%29%29%28x^2%2B1%29+%E2%88%92+2^x*2x%29%2F%28x^2%2B1%29^2+%2C+x+%3D+3

OpenStudy (anonymous):

Just click on Approximate Form

OpenStudy (anonymous):

so then is the answer 100 or was my other answer correct

OpenStudy (anonymous):

Wolfram Alpha is telling me that the answer is appox. 0.0745

OpenStudy (anonymous):

ok that is the answer that I got before

OpenStudy (anonymous):

p(x)=P(x)=2^x/(x^2+1) Diff. w .r.t. xType equation here. P’(x)=((x^2+1) 2^x log2-2^x*2x)/((〖x^2+1)〗^2 ) At x=3 P’(x)=((3^2+1)2^3 log2-2^3*2*3))/(3^2+1)^2 P’(x)=(10*8log2-8*6)/100 P’(x)=.8log2-.48